S 1, S 2 ,S 3 are the sum of n,2n.3n terms respectively of an A.P then a.S 3 =S 1 +S 2 b.S 3 =2(S 1 +S 2 ) c.S 3 =3(S 2 -S 1 ) d.non of these

S1,S2,S3 are the sum of n,2n.3n terms respectively of an A.P then
a.S3=S1+S2
b.S3=2(S1+S2)
c.S3=3(S2-S1)
d.non of these

Grade:12

1 Answers

Harsh Patodia IIT Roorkee
9 years ago
Hi Student
a= first term l=last term
S1= n/2(a+l)
S2=2n/2(a+l)= n(a+l)
S3= 3n/2(a+l) = n(a+l) + n/2(a+l) =S1+ S2

Clearly option A satisfies this.
….….….….….….….….…...